Postlab 7 3) Referring to the equipment of lab 7: a 10 g balli fired at 6 into a 280 g pendulum. What maximum angle will the pendulum swing to if the effective radius of the pendulum is 28 cm? Explain your solution.
Answer #1

10 g ball is ired into 280 g pendulum with 6 m/s

so they will move after collission with a combine speed let V

from momentum conservation

m1u1 + m2u2 = m1V + m2V

10 x 6 = ( 10 + 280 ) x V

V = 0.206 m/s

now

tan(theta)-(R-h)/R R-h u 0.206 m/s _ -10 280 0.29 kg

for h

from energy conservation

KEi + PEi = KEf + PEf

1/2 m V2 + 0 = 0 + mgh

0.5 x (0.206)2 = 9.81 x h

h = 0.22 cm

R = 28 cm

cos\theta = (28 - 0.22) /28

\theta = 7.2 degree answer